OpenStudy (anonymous):

When solving the system by the elimination method, what occurs when the second row is multiplied by 6 and the two equations are added?

OpenStudy (anonymous):

\[4x + 6y = -2\] \[3x - y = 9\]

OpenStudy (precal):

the y's will cancel out +6y-6y=0 then it will be easier to solve for x. We call this the elimination method. Once you know x, plug that into one of the original equations and solve for y.